Solve (d ^ 2 * y)/(d * x ^ 2) - 2tan x * (dy)/(dx) + 5y =


1
Expert's answer
2022-01-27T14:03:19-0500

We are given


"y''-2\tan x(y')+5y=e^x \sec x"

"\cos xy''-2\sin x(y')+5y\cos x=e^x"

We will be looking for a solution in the form


"u(x)=y(x)\cos x"

Then


"u'=y'\cos x-y\sin x"

"u''=y''\cos x-2y'\sin x-y\cos x"

Hence


"\cos xy''-2\sin x(y')=u''+u"

Substitute


"u''+u+5u=e^x"

Corresponding homogeneous differentional equation


"u''+6u=0"

Characteristic (auxiliary) equation


"r^2+6=0"

"r=\pm i\sqrt{6}"

The general solution of the homogeneous differentional equation is


"u_h=c_1\cos (\sqrt{6}x)+c_2 \sin (\sqrt{6}x)"

Find the partial solution in the form


"u_p=Ae^x"

"u_p'=Ae^x"

"u_p''=Ae^x"

Substitute


"Ae^x+6Ae^x=e^x"

"A=\dfrac{1}{7}"

"u_p=\dfrac{1}{7}e^x"

The general solution of the non homogeneous differentional equation is


"u=c_1\cos (\sqrt{6}x)+c_2 \sin (\sqrt{6}x)+\dfrac{1}{7}e^x"

Therefore the solution of the given differential equation is


"y=\dfrac{1}{\cos x}(c_1\cos (\sqrt{6}x)+c_2 \sin (\sqrt{6}x)+\dfrac{1}{7}e^x)"